OpenStudy (nincompoop):

find the formula for the sum of n terms then use the formula to find the limit as n -> infinity

OpenStudy (nincompoop):

\[\lim_{n \rightarrow \infty} \sum_{i=1}^{n}\frac{ 1 }{ n^3 }(i-1)^2\] setting (i-1) to u, so this would then be \[\lim_{n \rightarrow \infty} \sum_{i=1-1}^{n-1}\frac{ 1 }{ n^3 }(u)^2\]

OpenStudy (nincompoop):

WHAT THE HICKIE IS INFINITY - 1?

OpenStudy (lastdaywork):

Infinity - (real number) = infinity

OpenStudy (nincompoop):

@sourwing did I set the sigma correctly?

OpenStudy (nincompoop):

or instead of i = 1 it would be u = 1-1 \[\lim_{n \rightarrow \infty} \sum_{u=1}^{n}\frac{ 1 }{ n^3 }(i-1)^2\]

OpenStudy (nincompoop):

\[\lim_{n \rightarrow \infty} \sum_{u=1-1}^{n-1}\frac{ 1 }{ n^3 }(i-1)^2\]

OpenStudy (anonymous):

well, again the increment is i, not n. So n is like a constant

OpenStudy (nincompoop):

\[\lim_{n \rightarrow \infty} \sum_{u=1-1}^{n-1}\frac{ 1 }{ n^3 }(u)^2\]

OpenStudy (nincompoop):

so it's not even correct to touch n really?

ganeshie8 (ganeshie8):

pull n^3 out of sigma, the sum is just sum of squares of first (n-1) natural numbers

ganeshie8 (ganeshie8):

\(\large \lim \limits_{n \rightarrow \infty} \sum \limits_{u=1-1}^{n-1}\frac{ 1 }{ n^3 }(u)^2 = \lim \limits_{n \rightarrow \infty} \frac{ 1 }{ n^3 }\sum \limits_{u=0}^{n-1}(u)^2\)

OpenStudy (nincompoop):

\[\lim_{n \rightarrow \infty} \frac{ 1 }{ n^3 }\sum_{i=1}^{\infty }(i-1)^2\]

OpenStudy (anonymous):

if you let u = i-1, when i = 1, u = 0 when i = n, u = n-1 so, \[\lim n \rightarrow \infty \sum_{u=0}^{n-1} u^2\] which is the same as lim n rightarrow infty sum_{u=1}^{n-1} u^2 \[(1/6) u(u+1)(2u+1)\]

ganeshie8 (ganeshie8):

yes both are same, u or i... dont mess wid limit yet... .

OpenStudy (lastdaywork):

@nincompoop Forget the limit and solve sigma first. You can apply limit later.

OpenStudy (anonymous):

plug in n-1 for u (1/6) (n-1) (n-1+1) (2(n-1)+1)

OpenStudy (nincompoop):

\[\lim_{n \rightarrow \infty} \frac{ 1 }{ n^3 }\sum_{u=0}^{n-1 }(u)^2\]

OpenStudy (anonymous):

(1/6) (n-1)(n-1+1) (2(n-1)+1) / n^3 has limit of 1/2 as n->inf

OpenStudy (nincompoop):

I see what you did there @sourwing

OpenStudy (nincompoop):

you plugged -1 on all n's then proceeded with summation formulas

OpenStudy (anonymous):

yep

OpenStudy (anonymous):

Did i say 1/2? I meant 1/3 :DD

OpenStudy (nincompoop):

I was hoping you would correct that because I remember you said it's the fraction value that it approaches or something of that sort

OpenStudy (anonymous):

yeah ^^

OpenStudy (nincompoop):

I don't even need to evaluate different values that saves a lot of time